OpenStudy (anonymous):

Am i correct? Simplify i^41. i <--- my answer –1 –i 1

OpenStudy (jdoe0001):

hmm what's \(\bf i^4=?\) anyway?

OpenStudy (anonymous):

Yes, you are correct. It would be i. Since, you have 41 you would find a number that is lower and divides by the number 4. For example, in this case 40 divided by 4 is 10 and so that number would be 1 which is i^4 but your looking for 41so you start at the top again with i and go down but for this case it would only go down one and that would be i.

OpenStudy (anonymous):

\[\iota ^{41}=\iota ^{40}\iota =\left( \iota^2 \right)^{20}\iota=\left( -1 \right)^{20}\iota=?\]